OpenStudy (loser66):

Show that \(1^p + 2^p +\cdots + (p-1)^p \equiv 0(mod p)\) Please, help

OpenStudy (loser66):

I am looking for several methods. One of them is p > 1, 2, ...., (p-1) , and p is prime, hence (p, 1) = 1, (p, 2) =1.... and so one that is by Wilson's theorem, \(1^ p \equiv 1 (modp)\\2^p \equiv 2 (mod p)\\\cdots\\(p-1)^p \equiv p-1 (mod p)\) then, \(1^p +2^p +\cdots +(p-1)^p \equiv 1+2 +\cdots +(p-1) (mod p)\) But the right hand side is \(\sum_{i =1}^{p-1} i = \dfrac{(p-1)(p-1+1)}{2}= \dfrac{p(p-1)}{2}\) and p is odd, hence 2| p-1, then \(\dfrac{p(p-1)}{2}\equiv 0(modp)\)

OpenStudy (empty):

I'm kinda thinking we might be able to pair up the last and first terms: \[[1^p+(p-1)^p] + [2^p+(p-2)^p] + \cdots \equiv 0 \mod p\] Since distributing the binomials, and p is odd we will get: \[(p-n)^p \equiv -n^p \mod p\] so all the terms cancel each other out. I think if we wrote it out in summation notation it'd be more rigorous and also we could prove for p=2 independently since it's the only even prime.

OpenStudy (loser66):

Yes, it works to me. :) Thank you

OpenStudy (empty):

Well I already typed this up, so here we go I guess haha. I guess since \(1 \cancel \equiv 0 \mod 2\) is false, p is a prime greater than 2 for this problem. Oh well, no big deal. \[\sum_{n=1}^{p-1} n^p \mod p\] \[\sum_{n=1}^{(p-1)/2} n^p + (p-n)^p \equiv 0 \mod p\]

OpenStudy (loser66):

Again, thank you so much.